I struggle with these kind of questions as well.. I apprciate any help.


3/5(-15x-20)=21


Thank you!

Answer by mananth(16946)

3/5(-15x-20)=21
when there is a multiplication of fraction eliminate the denominator
multiply by 5
3/5 * 5 (-15x-20)=21*5
3(-15x-20)=105
divide by 3
(-15x-20)= 35
add 20
-15x=35+20
-15x=55
x=55/-15
x=-11/3
...
m.ananth@hotmail.ca